The force between two stationary charged objects is measure to be 15.0 N. If distance between the objects is doubled, what is the new force between the two objects?
1
Expert's answer
2019-04-15T10:06:55-0400

The force between two stationary charged objects is 


"F=\frac{kq_1q_2}{r^2}"

The new force:


"F=\frac{kq_1q_2}{(2r)^2}=\frac{kq_1q_2}{4r^2}=\frac{F}{4}=\frac{15}{4}=3.75 N"